### 22 Mar 2026 v1

# KG-Hopper: Empowering Compact Open LLMs with Knowledge Graph Reasoning via Reinforcement Learning

Abstract
Large Language Models (LLMs) demonstrate impressive natural language capabilities but often struggle with knowledge-intensive reasoning tasks. Knowledge Base Question Answering (KBQA), which leverages structured Knowledge Graphs (KGs) exemplifies this challenge due to the need for accurate multi-hop reasoning. Existing approaches typically perform sequential reasoning steps guided by predefined pipelines, restricting flexibility and causing error cascades due to isolated reasoning at each step. To address these limitations, we propose KG-Hopper, a novel Reinforcement Learning (RL) framework that empowers compact open LLMs with the ability to perform integrated multi-hop KG reasoning within a single inference round. Rather than reasoning step-by-step, we train a Reasoning LLM that embeds the entire KG traversal and decision process into a unified “thinking” stage, enabling global reasoning over cross-step dependencies and dynamic path exploration with backtracking. Experimental results on eight KG reasoning benchmarks show that KG-Hopper, based on a 7B-parameter LLM, consistently outperforms larger multi-step systems (up to 70B) and achieves competitive performance with proprietary models such as GPT-3.5-Turbo and GPT-4o-mini, while remaining compact, open, and data-efficient. The code is publicly available at: https://github.com/Wangshuaiia/KG-Hopper.
I Introduction
Large Language Models (LLMs) have achieved remarkable success across diverse domains but still exhibit notable shortcomings, such as hallucinations and factual inaccuracies, particularly in knowledge-intensive tasks [wang2022knowledge]. This is largely due to the implicit storage of knowledge within their model weights, making knowledge updates cumbersome and resource-intensive through fine-tuning [wang2025domain]. To address this limitation, Retrieval-Augmented Generation (RAG) has emerged as an effective strategy, enabling LLMs to dynamically access external knowledge during inference [wang2025plugging]. Among various external resources, Knowledge Graphs (KGs) stand out as structured and reliable knowledge bases, offering explicit, interpretable, and easily updateable knowledge beneficial in critical applications such as medicine and finance [wang2024knowledge].
Knowledge Base Question Answering (KBQA), which aims to leverage structured knowledge from KGs to answer questions, often requires complex multi-hop reasoning—traversing multiple interconnected relationships within a KG [wang-yu-2025-iquest]. Current approaches typically follow step-by-step reasoning strategies, sequentially processing entities and relationships from a predefined pipeline. For example, as illustrated in Figure 1(a), answering the question "What is the official flower of the area affected by Tropical Storm Fabio?" involves first identifying the affected area and subsequently retrieving its official flower. However, such rigid frameworks present critical drawbacks:
Figure 1: Multi-step vs one-round multi-hop reasoning over a knowledge graph: (a) multi-step reasoning and (b) our one-round reasoning. The multi-step pipeline invokes multiple sequential LLM calls and fails due to the missing entity Yellow Hibiscus, leading to an incorrect path. In contrast, our one-round approach performs the entire reasoning process within a single Reasoning LLM call, maintaining coherence and demonstrating robustness to incomplete knowledge.
  • Limited Flexibility and Local Optima Susceptibility. Traditional methods sequentially aggregate information through individual inference steps guided by predefined reasoning paths. Consequently, these methods struggle to dynamically adjust when facing incomplete or misleading KG data. For instance, as shown in Figure 1(a), missing the critical entity "Yellow Hibiscus" could mislead a multi-step method employing beam search into incorrectly choosing the national flower of Mexico, with limited capability for backtracking.
  • Error Cascading and Reasoning Bias. Stepwise methods inherently propagate errors from earlier steps. Incorrectly choosing a node, such as "Mexico" in Figure 1(a), directly influences subsequent reasoning steps. Additionally, treating each reasoning step independently neglects inter-step dependencies, causing biases and potential deviations from the original query intent.
To mitigate these limitations, two lines of research offer promising advances. First, Reinforcement learning (RL) have demonstrated its effectiveness in navigating discrete and combinatorial decision spaces, such as KG traversal by optimizing reasoning policies through exploration and long-term rewards [xiong2017deeppath, das2018go, lin2018multi]. Second, recent work on Reasoning LLMs, such as ChatGPT-o1 [jaech2024openai] and DeepSeek-R1 [guo2025deepseek], has significantly improved the reasoning capabilities of language models. These models perform a dedicated “thinking” phase before generating answers, where each token in the thought sequence can be viewed as a latent variable. Importantly, later tokens can revise earlier ones, enabling self-reflection and correction within the reasoning process [zhu2025chain]. This iterative refinement mechanism aligns naturally with the multi-hop nature of complex KBQA tasks, allowing LLMs to reason more coherently and correct potential errors before answer generation.
Building on these insights, we propose KG-Hopper, a novel KBQA framework integrating the entire multi-hop reasoning process within a single LLM call, as depicted in Figure 1(b). Specifically, KG-Hopper leverages RL to enhance the reasoning capabilities of LLMs by training them to utilize KG retrieval tools effectively.
KG-Hopper addresses previous limitations with the following advantages: (1) By embedding the complete reasoning chain within a single inference step, KG-Hopper captures cross-step dependencies, mitigating local biases and ensuring coherent reasoning. (2) Integrating reasoning into the Reasoning LLM’s "thinking" phase allows flexible exploration and effective backtracking, significantly reducing cascading errors without predefined reasoning patterns. (3) Our framework leverages RL to explicitly guide and refine reasoning processes, substantially improving the model’s performance in complex multi-hop scenarios. (4) It remains compact and deployable, using a 7B LLM to achieve competitive results with much larger models.
To effectively train KG-Hopper, we initially generate cold-start data and employ Supervised Fine-Tuning (SFT), enabling autonomous KG retrieval. We subsequently strengthen multi-hop reasoning capabilities through targeted RL training, guided by carefully designed reward functions for retrieving and reasoning process. Additionally, we mask the retrieved triples from KG to minimize external knowledge interference and employ history resampling to enhance training efficiency.
Our main contributions are as follows:
  • We introduce a RL framework for LLM–KG integration, combining structured reward signals for retrieval and reasoning with masked supervision over KG content. To the best of our knowledge, this is the first work to apply RL to enable end-to-end KG reasoning within LLMs.
  • We propose a one-round KBQA framework that performs multi-hop reasoning entirely within the “thinking” phase of a single LLM inference, enabling autonomous and iterative KG traversal without multi-step orchestration.
  • We achieve strong empirical results on eight KBQA benchmarks: our compact 7B-parameter model consistently outperforms multi-step methods using models up to 70B, and matches or exceeds the performance of proprietary models such as GPT-3.5-Turbo and GPT-4o-mini.
II Task Definition
We address the task of Knowledge Base Question Answering (KBQA) over a knowledge graph $G=\{(e,r,e^{\prime})\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\}$, where each triple $(e,r,e^{\prime})$ denotes a relation $r$ between two entities $e$ and $e^{\prime}$. Given a natural language question $q$, along with an identified topic entity $e_{t}\in\mathcal{E}$ mentioned in the question, the objective is to infer the correct answer entity $e_{a}\in\mathcal{E}$ by reasoning over the knowledge graph $G$. The answering process typically starts from the topic entity $e_{t}$ and involves exploring the graph to identify the answer entity $e_{a}$. In practice, large-scale knowledge graphs are often sparse and incomplete, making it difficult to locate the answer through simple one-hop queries. As a result, multi-hop reasoning is frequently required, where the answer entity $e_{a}$ may reside several hops away from the topic entity $e_{t}$.
III Method
Our approach leverages reinforcement learning to train LLMs for KG retrieval and reasoning. This section introduces two key components: the construction of a knowledge graph retrieval tool, and the RL framework for training LLMs to interact with the knowledge graph.
III-A Knowledge Graph Retrieval Tool
Large-scale knowledge graphs are typically stored in graph databases and queried using SPARQL for retrieval. Our retrieval process starts from a topic entity and expands the search. Due to the vast number of triples connected to a single node, it is crucial to filter out irrelevant information. To address this, we adopt a two-stage strategy: first retrieving the set of directly connected edges (predicates) and then selecting the ones relevant to the input query.
Given a topic entity, our tool first retrieves all predicates and corresponding object entities linked to it. For example, given the question “What books did J.K. Rowling write?”, the topic entity is J.K. Rowling. The following SPARQL query is issued to retrieve all outgoing predicates and their objects:
SELECT ?predicate ?object
WHERE {
ns:m.05b6w ?predicate ?object .
Here, ns:m.05b6w denotes the Freebase ID for J.K. Rowling, ?predicate retrieves all relations where she appears as the subject, and ?object returns the corresponding object entities.
For example, the retrieved predicates include:
   ns:book.author.works_written,
   ns:people.person.place_of_birth,
   ns:people.person.nationality.
Based on the semantic relevance to the question, the most related predicate, such as ns:book.author.works_written, is selected. This predicate indicates the books authored by the subject. Next, the tool issues another SPARQL query to retrieve the tail entities (i.e., books) associated with the selected predicate:
SELECT ?tailEntity
WHERE {
ns:m.05b6w ns:book.author.works_written ?tailEntity .
This query fetches all book entities linked to J.K. Rowling through the works_written relation, effectively answering the question.
In summary, the retrieval tool takes an entity as input and returns a set of relevant triples from the knowledge graph by identifying and filtering meaningful relations and associated entities.
III-B Cold Start
To avoid the unstable cold-start phase typically seen in early RL training. Specifically, we construct and collect a small set of CoT annotated data to fine-tune the base LLM, which is then used as the initial RL actor. The cold-start dataset is created with two primary objectives: (1) to demonstrate how the model should properly invoke the knowledge graph retrieval tool; and (2) to enforce a consistent, structured format for answer generation.
We define a rule such that when the LLM generates the special tokens <search> and </search>, it triggers the KG retrieval tool to search the enclosed entity. The retrieved triples are then wrapped with <searched_triples> tags and appended to the current context, enabling the LLM to continue generation with access to the retrieved knowledge. To collect cold start data, we use few-shot prompting with a long CoT example to elicit responses from a powerful LLM. We then select examples that correctly invoke KG queries and exhibit high readability. For instance, given the question: what timezone is Utah in? A preferred response would begin with an explicit motivation for search, such as:
<think>Given the information, I don’t have specific knowledge about Utah’s timezone. I will perform a search for the required information. The query to search for is <search>utah</search> <searched_triples> (Utah, timeZone, Mountain Time Zone)</searched_triples>. Based on the search results, I found that Utah is in the Mountain Time Zone.</think> <answer>Mountain Standard Time</answer>
We select such examples, which exhibit clarity, appropriate tool invocation, and well-structured reasoning, as the cold-start training data. Such structure is the desired output format of the LLM. The <think> section encapsulates the full CoT reasoning process, including autonomous invocation of the KG tool (<search>), retrieval results (<triples>), and intermediate deductions. The <answer> section then summarizes the reasoning into a final answer. This also represents the desired output format of the LLM.
We use the collected data to fine-tune the base LLM. During training, we mask the tokens within the <triples> to prevent the model from being distracted by retrieved knowledge. This encourages the model to generalize its reasoning strategy while preserving tool-use behaviors. The fine-tuned model is then further optimized via reinforcement learning.
III-C Reasoning-oriented Reinforcement Learning
Given the difficulty of obtaining sufficient high-quality long CoT reasoning data, we leverage Reinforcement Learning with Human Feedback (RLHF) as a principled alternative to explicitly guide the model toward effective multi-hop reasoning. We design a composite reward function to guide the model’s behavior throughout the reasoning process. The reward comprises four components: retrieval reward, format reward, reasoning reward, and final answer reward.
  • Retrieval Reward To encourage the model to search for answers through the knowledge graph rather than relying solely on its internal knowledge, we provide a positive reward for each invocation of the query tool. However, to prevent the model from overusing the tool solely for reward accumulation, we apply a cap on the maximum reward. The retrieval reward is defined as:
    $R_{\text{search}}=\min(0.5\cdot n,0.8)$ (1)
    where $n$ denotes the number of times the query tool is invoked. This design incentivizes query usage while discouraging excessive or redundant retrievals.
  • Format Reward To enforce structured reasoning and tool usage, we define a format reward that requires the generated text to follow a predefined format. Specifically, the model must use the tags <think>, <search>, and <answer> appropriately. If all tags appear in the correct positions and order, a fixed reward is granted:
    $R_{\text{format}}=\begin{cases}0.5&\text{if the format meets all requirements}\\ 0&\text{otherwise}\end{cases}$ (2)
    This constraint ensures that the model explicitly separates its reasoning, search actions, and final answer.
  • Reasoning Reward In multi-hop KG reasoning, an error at any intermediate step can lead to an incorrect final answer. To mitigate this, we introduce a reward signal that directly evaluates the quality of the reasoning process itself, encouraging the model to make sound decisions at each step, rather than focusing solely on the final answer. The model is expected to adapt its behavior based on the informativeness of retrieved triples. When the retrieved information is sufficient, the model should organize and synthesize it into a coherent answer. Otherwise, it should continue reasoning by selecting the next entity or relation to query. In cases where the KG lacks the necessary facts (e.g., due to incompleteness), the model is allowed to fall back on its internal knowledge to infer a plausible answer. To assess the quality of the model’s reasoning behavior, we use an external LLM to evaluate the full reasoning trace enclosed in the <think> tag. The reasoning reward is computed as:
    $R_{\text{reason}}=f_{r}(\text{reasoning process})\in(0,1)$ (3)
    where $f_{r}$ denotes an external LLM, and a higher score indicates a more reasonable and logically valid reasoning trace.
  • Answer Reward Finally, we reward the model if it provides a correct answer. Since the generated answer may differ from the ground truth due to variations such as abbreviations or phrasing, we again use a separate LLM to perform semantic similarity assessment between the predicted answer and the ground truth (LLM as Evaluator). The final answer reward is defined as:
    $R_{\text{answer}}=f_{a}(\text{predicted\_answer},\text{ground\_truth})\in\{0,1\}$ (4)
    where $f_{a}$ denotes an external LLM, and the predicted answer is extracted from the <answer> tag. A reward of 1 is given if the LLM determines the answer is semantically correct, and 0 otherwise. In the absence of ground truth, we directly use a (preferably larger) LLM with rich knowledge as both the Judge and Evaluator, i.e., we rely on the LLM’s internal knowledge to assess whether the predicted answer is correct.
By combining the above four reward functions, the total reward provides comprehensive guidance to the model throughout the reasoning process, promoting structured, accurate, and knowledge-grounded answers, which can be represented as:
$R_{\text{final}}=R_{\text{search}}+R_{\text{format}}+R_{\text{reason}}+R_{\text{answer}}$ (5)
III-D Optimization
We optimize the reasoning policy using Group Relative Policy Optimization (GRPO), which trains the LLM to maximize the expected reward of generated reasoning trajectories. Formally, the objective is simplified as:
$\theta^{*}=\arg\max_{\theta}\;\mathbb{E}_{q\sim\pi_{\theta}}\big[R_{\text{final}}(q)\big],$ (6)
where $\pi_{\theta}$ denotes the LLM policy and $R_{\text{final}}(q)$ is the corresponding reward function. GRPO estimates relative advantages within sampled output groups to stabilize optimization and encourage higher-quality reasoning paths.
Masking Retrieved Triples. Retrieved triples are provided as auxiliary context but are not expected to be generated by the model. We therefore mask tokens enclosed by <triples> and </triples> during loss computation to prevent the model from learning to reproduce retrieved content.
History Resampling for Efficient Training. In KBQA, simple one-hop queries often produce uniformly high rewards, leading to near-zero normalized advantages and inefficient learning. Following a history resampling strategy [zhang2025srpo], we remove one-hop questions after an initial training phase, encouraging the model to focus on multi-hop reasoning in a curriculum learning manner [narvekar2020curriculum].
IV Experiments
IV-A Datasets
We evaluate our approach on eight widely-used datasets for KBQA, leveraging two large-scale general-purpose knowledge graphs: Freebase and WikiData. Four of the datasets are based on the Freebase knowledge graph and are standard KBQA benchmarks: ComplexWebQuestions (CWQ) [talmor2018web], WebQuestionsSP (WebQSP) [yih2016value], WebQuestions [berant2013semantic], and GrailQA [gu2021beyond]. The other four datasets are grounded in the WikiData knowledge graph. Among them, QALD10-en [perevalov2022qald] is a KBQA dataset, while T-REx [elsahar2018t] and Zero-Shot RE [petroni2021kilt] are designed for slot filling tasks, and Creak [onoe2creak] focuses on factual verification. Following prior work [sun2024thinkongraph, zhao2024kg, xiong-etal-2024-interactive], we use Hit@1 score as the evaluation metric for both question answering and slot filling tasks, while accuracy is used to evaluate performance on the Creak dataset.
IV-B Implementation Details
We conduct experiments using two instruction-tuned language models as backbones: LLaMA-3.1-8B-Instruct and Qwen-2.5-7B. To mitigate the unstable cold-start phase of reinforcement learning, we first construct a set of 500 high-quality examples to teach the models how to properly invoke the knowledge retrieval tool. We randomly sample 2,000 examples from 8 datasets for RL training. The scoring model $f_{r}$ for the reasoning process is Llama-3.3-70B, while the model $f_{a}$ used to determine whether the predicted answer matches the ground truth is Llama-3.2-3B. Starting from the second epoch, we apply a resampling strategy to dynamically filter out trivial questions and retain more informative ones for continued training. The training is performed on 8 NVIDIA A100 80G GPUs in total. For each input query, we generate 16 outputs (rollouts). We train for 2 epochs with a batch size of 16 and a learning rate of $1\text{e}{-6}$. The rollout temperature is set to 1, the PPO clip ratio is 0.2, and the KL divergence penalty coefficient is $1\text{e}{-5}$.
IV-C Main Results
We design our experiments to explore the reasoning effectiveness, training efficiency, and design trade-offs of our method KG-Hopper. KG-Hopper is designed as an open-source, compact, and flexible solution that promotes transparency in the training process and supports targeted modifications. Through our experiments, we investigate how RL contributes to efficient multi-hop KBQA. Results are reported in Table LABEL:tab:grouped-llm-comparison.
In terms of integration strategy, prompt-only LLMs, including strong proprietary models like GPT-4o, consistently underperform on multi-hop KBQA tasks. This highlights the limitations of relying solely on parametric knowledge and the lack of explicit, structured reasoning. Adding KG retrieval tools improves performance significantly (often by 10–30 Hits@1), but without any form of model-level adaptation, these systems exhibit fixed reasoning patterns and struggle to generalize beyond shallow queries. SFT on KG reasoning tasks further improves results, especially for moderately complex questions. However, its imitation-based learning process tends to be brittle, it teaches the model to reproduce specific reasoning paths rather than adaptively exploring alternatives. Our RL-trained model, KG-Hopper, consistently outperforms SFT-based models of similar (7-8B) or larger size (13B, 70B) and matches or exceeds the performance of GPT-4o-mini + KG, particularly on more complex multi-hop reasoning tasks.
TABLE I: Performance comparison (Hits@1). Results marked with ‘*’ are taken directly from the corresponding original papers. Bold numbers indicate the best performance, while underlined numbers denote the second-best.